express in exponential form: log squared 1/8=-3
sorry that i couldnt find the squared symbol so i just wrote it in. thanks and i hope you can help me with this question

Answer by stanbon(75887)

log squared 1/8=-3
"log squared" sound mighty peculiar.
Are you sure it's not (base 2)
or maybe the (1/8) is squared.
I think it is probably base2.
Assuming it is the problem goes like this:
log(base2)(1/8)=-3
2^(-3)=1/8
2^(-3)=2^(-3)
A mysterious problem. I suggest you check it out again.
Cheers,
Stan H.